OpenStudy (gschibby):

If I have a function f(x) = x^2 * e^-x and a line L that tangents the graph of the function at x =1. How do I find the function for the line L ?

hartnn (hartnn):

find the slope of line L by differentiating f(x) slope = f'(x) at x=1 = f'(1) can you find f'(1) first ?

hartnn (hartnn):

once you have slope, you will just need a point on L, when x=1, find f(1), and the point will be (1, f(1)) try it, if you get stuck , i am here :)

OpenStudy (gschibby):

so \[f'(x) = 2xe ^{-x}-x ^{2}e ^{-x}\]

hartnn (hartnn):

yes, put x=1

OpenStudy (gschibby):

I get 0,3678794412 :P

OpenStudy (gschibby):

or \[e ^{-1}\]

hartnn (hartnn):

keep it as 1/e :)

OpenStudy (gschibby):

so the point is \[(1,e ^{-1})\]

hartnn (hartnn):

yes

hartnn (hartnn):

we have slope, m= 1/e, point = (1,1/e) just use slope point form y-y1 = m (x-x1)

hartnn (hartnn):

could you find the function for line ?

hartnn (hartnn):

there?

OpenStudy (gschibby):

Just looking at ProfRObBob on youtube for the slope-stuff :P I'm norwegian so the english terms are a little different than ours ;)

hartnn (hartnn):

ohh

hartnn (hartnn):

y-1/e = 1/e (x-1) y -1/e = x/e -1/e y = x/e or x =ey is the function for line L hope you got the same :)

OpenStudy (gschibby):

Yup, thanks a lot! :)

hartnn (hartnn):

welcome ^_^